ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 20G25 Abels, Herbert, Universitaet Bielefeld, Germany Reductive groups as metric spaces I would like to report on work in progress jointly with G. Margulis on the following topics:

1) A characterization of the metrics whose restriction to a maximal split torus is a norm,

2) a question of Siegel where he compares the metric in the fundamental domain for an arithmetc group Γ with that of the orbit space modulo Γ, and

3) undistortedness of Γ.